OpenStudy (anonymous):

Stan invested $17,000, pat at 18% and part at 4%. If the total interest at the end of the year is $1,380, how much did he invest at each rate?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

if he invested P at 18% then he invested the rest, 1700-P and 4% set \[.18P+.04(1700-P)=1380\] and solve for P

OpenStudy (anonymous):

\[.18P+68-.04P=1380\] \[.12P=1380-680=1312\] \[P=\frac{700}{.12}=583\tfrac{1}{3}\]

OpenStudy (anonymous):

typo on first line. should be \[.18P+680-.04P=1380\]

OpenStudy (anonymous):

damn and on the second line too! should be \[.12P=1380-680=700\] answer is right tho...

OpenStudy (anonymous):

so the answer is 700 than right